OpenStudy (anonymous):

please help substitution 2x-3y=8 x+y=11

OpenStudy (cwrw238):

from second equation x = 11 - y plug this into the first equation and solve for y then plug this value of y into second equation to find value of x

OpenStudy (anonymous):

so 2(11-y)3y=8

OpenStudy (princeharryyy):

these equations don't need substitution just multiply equation 2 by 2 and subtract from 1 to get the answer @ivonlorena

OpenStudy (princeharryyy):

-5y = -14 =>y = 14/5 solve for x after that @ivonlorena

OpenStudy (princeharryyy):

x=11-14/5

OpenStudy (anonymous):

so x=11 and y=14/5

OpenStudy (princeharryyy):

x=41/5

OpenStudy (anonymous):

so whats y

OpenStudy (princeharryyy):

y=14/5